# Signal Processing Meets SGD: From Momentum to Filter

### v5

@@ 4.1 Empirical Evaluation
Hyperparameter tuning. We conducted careful hyperparameter tuning [66] in our experiments. For image classification, we used the following settings:
  • SGDF: We used the default parameters of Adam: $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$.
  • SGD: We set the momentum as 0.9, which is the default for many networks such as ResNet and DenseNet. We searched for the learning rate among {10.0, 1.0, 0.1, 0.01, 0.001}.
  • Adam, RAdam, MSVAG, AdaBound: We searched for the optimal $\beta_{1}$ among $\{0.5,0.6,0.7,0.8,0.9\}$, searched for $\alpha$ as in SGD, and set other parameters to their default values in the literature.
  • AdamW, SophiaG, Lion: We used the same parameter searching scheme as Adam. For other optimizers, we set the weight decay to $5\times 10^{-4}$; for AdamW, since the optimal weight decay is typically larger [38], we search weight decay among $\left\{10^{-4},5\times 10^{-4},10^{-3},10^{-2},10^{-1}\right\}$.
  • SophiaG, Lion: We searched for the learning rate among$\{10^{-3},10^{-4},10^{-5}\}$ and scaled Lion’s learning rate [36]. According to [36, 10], we set $\beta_{1}$=0.965, 0.9 and $\beta_{2}$=0.99 as default parameters.
@@
Image Generation. Stability of optimizers is crucial in practical applications such as training Generative Adversarial Networks (GANs). GAN training involves alternating between the generator and discriminator in a mini-max game and is typically unstable; SGD often leads to model collapse, while adaptive methods such as Adam and RMSProp are recommended in practice. Therefore, training GANs serves as a good test for stability. To facilitate result reproduction, we provide the parameter table for this subpart in Tab. 7.
@@
We conducted experiments with one of the most widely used models, the Wasserstein-GAN with gradient penalty (WGAN-GP) [47], using a small model with a vanilla CNN generator. Employing popular optimizers [4, 61], we trained the model for 100 epochs, generating 64,000 fake images from noise and computed the Frechet Inception Distance (FID) [24] between the fake images and the real dataset (consisting of 60,000 real images). The FID score captures both the quality and diversity of generated images and is widely used to assess generative models (lower FID is better). For SGD and MSVAG, we used results from [66]. We conducted 5 runs of experiments and reported the results in Fig. 3.

### v6

@@ 4.1 Empirical Evaluation
Hyperparameter tuning. Following Zhuang et al. [66], we delved deep into the optimal hyperparameter settings for our experiments. In the image classification task, we employed these settings:
  • SGDF: We adhered to Adam’s original parameter values: $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$.
  • SGD: We set the momentum to 0.9, the default for networks like ResNet and DenseNet. The learning rate was searched in the set {10.0, 1.0, 0.1, 0.01, 0.001}.
  • Adam, RAdam, MSVAG, AdaBound: Traversing the hyperparameter landscape, we scoured $\beta_{1}$ values in $\{0.5,0.6,0.7,0.8,0.9\}$, probed $\alpha$ as in SGD, while tethering other parameters to their literary defaults.
  • AdamW, SophiaG, Lion: Mirroring Adam’s parameter search schema, we fixed weight decay at $5\times 10^{-4}$; yet for AdamW, whose optimal decay often exceeds norms [38], we ranged weight decay over $\left\{10^{-4},5\times 10^{-4},10^{-3},10^{-2},10^{-1}\right\}$.
  • SophiaG, Lion: We searched for the learning rate among $\{10^{-3},10^{-4},10^{-5}\}$ and adjusted Lion’s learning rate [36]. Following  [36, 10], we set $\beta_{1}$=0.965, 0.9 and $\beta_{2}$=0.99 as the default parameters.
@@
Image Generation. The stability of optimizers is crucial, especially when training Generative Adversarial Networks (GANs). If the generator and discriminator have mismatched complexities, it can lead to imbalance during GAN training, causing the GAN to fail to converge. This is known as model collapse. For instance, Vanilla SGD frequently causes model collapse, making adaptive optimizers like Adam and RMSProp the preferred choice. Therefore, GAN training provides a good benchmark for assessing optimizer stability. For reproducibility details, please refer to the parameter table in Table 7.
@@
We evaluated the Wasserstein-GAN with gradient penalty (WGAN-GP) [47]. Using well-known optimizers [4, 61], the model was trained for 100 epochs. We then calculated the Frechet Inception Distance (FID) [24] which is a metric that measures the similarity between the real image and the generated image distribution and is used to assess the quality of the generated model (lower FID indicates superior performance). Five random runs were conducted, and the outcomes are presented in Figure 3. Results for SGD and MSVAG were extracted from Zhuang et al. [66].